Keisha is 12 years old. Her mother is (m) years older than she.
How old will her mother be in 5 years' time? Give the answer in terms of m.
If m=20 how old will her mother be in five years time?

Answer by richwmiller(17219)

k=12
r =k+m
r is mothers age now.
(r+5)=k+m+5 in five years
r+5=(12+20)+5
r+5=37 in 5 years